OpenStudy (anonymous):

If (x-3)(x+3)=a, then (2x-6)(x+3)=? 2a 3a 4a 6a 9a

OpenStudy (anonymous):

If you simplify (2x-6) = 2(x-3) 2 is a common factor, I think you can proceed from there

OpenStudy (anonymous):

I already found that out, but I'm still confused as to what to do about it @Florcelda

OpenStudy (anonymous):

(2x-6)(x+3) = 2(x-3)(x+3) but (x-3)(x+3) = a right?

OpenStudy (anonymous):

So 2(x-3)(x+3) = 2a